# CaloClouds3: Ultra-Fast Geometry-Independent Highly-Granular Calorimeter Simulation

### v1

@@ 1 Introduction and Motivation
In this paper, we present the latest of a series of models designed for the fast simulation of photon showers in a high granularity calorimeter. This is the latest of the CaloClouds series of models, which represent the showers to be modeled as pointclouds. Two ML techniques are used; a normalising flow combined with a diffusion model. For a review covering other possible architectures, see [43].
@@
The previous model, CaloClouds2, was only capable of simulating photons incident at a single angle and position. Incident angle conditioning, along with location agnostic training data, allows a single model to be used for all photons entering the calorimeter, which simplifies training, and potentially occupies less memory at inference time. The octagonal barrel shape and magnetic field, which is aligned with the beam axis, break the rotational symmetry about the shower axis, therefore the model must be conditioned on both $\theta$ and $\phi$ directions separately.
@@
In section 2 we describe the example detector we have chosen to demonstrate our architecture on, and the dataset generation is detailed. In section 3 the new CaloClouds3 model is defined. In sections 4 and 5 we compare physics accuracy between the previous iteration, CaloClouds2, and the new model, CaloClouds3. After this, a snapshot of the newly attained capacity for full reconstruction is offered in section 5, however this is covered in more depth in a companion paper. In the penultimate section, section 6, the improvements in inference speed between CaloClouds2 and CaloClouds3 are demonstrated. Finally, we draw the conclusions in section 7.
@@ 2.2 Generation and preprocessing
Photons between $1$ and $127~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ were generated using Geant4, a slightly expanded range from the energies considered in CaloClouds2. In the framework used here, a fast model is only utilised (triggered) once a photon in an appropriate energy range reaches the surface of the calorimeter, therefore any photons that start showering before the surface of the calorimeter will not be simulated with the fast simulation model [^2] Potentially, some secondary photons from the shower of an early showering photon could reach the calorimeter surface with sufficient energy to trigger the fast simulation, but this would anyway have the same physics as a low energy primary photon, and so is covered by the dataset. therefore it is important to generate the datasets used to train the model without any pollution for early showering photons [^3] For further description of how the model is triggered in the context of a complete event simulation, see section 3.3.. To avoid any early showering, the photons in the training set are inserted into the simulation at the surface of the calorimeter. The angle of each photon is consistent with having originated from the Interaction Point (IP), having $30<\theta<150$ and $80<\phi<110$, sampled to have an even density on the upper barrel segment. Geant4 then records the hits of the shower in the sensitive layers. The training data includes $3\times 10^{6}$ events, with a test set of $7\times 10^{4}$ which was held back.
@@
The centre of each layer is shifted in the $x$ and $y$ directions to remove the shower’s tilt. This shift corresponds to the rotation that would align the incident photon with the $z$ axis, so the size of the shift is calculated from the depth of the layer $d_{\mathrm{layer}}$ the momentum of the incident particle $\vec{p}_{\mathrm{inc}}$; in the $x$ direction, the shift is $\Delta x=d_{\mathrm{layer}}\tan(\sin^{-1}(\vec{p}_{\mathrm{inc},x}/|\vec{p}_{\mathrm{inc}}|))$, and likewise for the $y$ direction. To first approximation, the shifted showers are now a similar oblong shape, somewhat akin to one created by an upward moving particle. This preprocessing helps the diffusion model to learn the details of backscattering, rather than being forced to focus on getting the gross shape right, which is known from the direction of the incident particle.
Finally, all 30 layers of the detector are trimmed to a box of $100\times 100$ cells, cutting to $\pm 250~$\mathrm{m}\mathrm{m}$\$ about the shower axis. To strike a balance between the fidelity of all Geant4 hits, and computational efficiency, the energy in each layer is binned into a grid of $500\times 500$ bins (each cell is split into $5\times 5$ bins). This is a slightly coarser granularity than that of CaloClouds2, and was implemented because lowering the number of points was seen to have minimal impact on the kinematics after projection, while improving the computational efficiency of the model. As with CaloClouds2, in each event, this grid is given a uniform random offset of up to half a cell, to reduce discretisation effects. A point with the total energy contained in the bin is created at the location of the highest energy point in the bin. This creates a set of up to $6,000$ points per shower.
@@ 3.1 Previous CaloClouds2 model
The key points for the following discussion are that the CaloClouds models are a composed of a normalising flow, referred to as ShowerFlow, which determines the overall structure of each photon shower, and a diffusion model, which generates each point. Both models are conditioned to create behaviour appropriate for the incident photon. The previous CaloClouds2 model was only capable of simulating photons at a fixed position, and the model’s hyperparameters were adopted from standard practice for these models.
@@ 3.2 Upgrade methodology
Conditioning on the incident angle is a relatively simple change. In CaloClouds2 the normalising flow was conditioned on $E_{\mathrm{inc}}$, the incident particle energy, in CaloClouds3 it will be conditioned on $(E_{inc},\vec{p}_{inc}/|\vec{p}_{inc}|)$ , the incident energy and normalised momentum direction. Likewise, the diffusion model of CaloClouds3 will be conditioned on $(E_{inc},\vec{p}_{inc}/|\vec{p}_{inc}|)$. Conditioning the diffusion model on the number of points in the shower, used in CaloClouds2, was found to not have significant effect and therefore removed.
@@
In order to limit the computational resources (and therefore the carbon footprint) required to investigate variations of normalising flow model, some toy models, representing expected data characteristics were employed. These small scale experiments focused on investigating two potential properties of a probability distribution function (pdf); sub manifolds and first or second order discontinuities. Physics data might be expected to contain sub manifolds, as the number of output parameters generally exceeds the input parameters, and would be expected to be approximately smooth, with few first order discontinuities. For example, the pdf of the energy in each layer of the detector would be expected to vary smoothly with incident particle energy, and have an approximate submanifold defined by adjacent layers having similar energies. Toy datasets were created from various pdfs, with and without the aforementioned properties. By observing the training and goodness of fit for normalising flow models with a similar architecture on those toy datasets, we could economically obtain indications of what might be expected in more complex cases. While this is far from conclusive, it helped guide development of the more expensive training process of the full model.
All expected characteristics of the physics data favoured the use of simple normalising flow models. Reducing the parameter space of our ShowerFlow model, it was seen that the number of parameters could be safely reduced by a factor of 5 (but that a factor of 10 was detrimental to performance). In a similar fashion, the number of parameters in the diffusion model was reduced by a factor of 4.
@@
While the toy models showed no overwhelming failure to learn due to manifold overfitting, submanifolds did somewhat impair training stability. Identifying and removing submanifolds in the normalising flow of our physical model still offered a marginal performance increase. Comparing the training process of ShowerFlow in CaloClouds2 and CaloClouds3 (figure 2), there are indications that a more stable training is taking place in CaloClouds3.
@@ 3.3 CaloClouds3 architecture
Figure 3: The model architecture of CaloClouds3. Parts belonging to ShowerFlow are in purple, and those belonging to the diffusion model are in green. A full description is provided in section $3.$
As with CaloClouds2, the model is composed of a normalising flow, referred to as ShowerFlow, (shown in purple in figure 3) and a distilled diffusion model (shown in green in figure 3). ShowerFlow is now limited to distributions along the shower axis, specifically the points and energy per layer, and the diffusion model is still used to create individual points. Both are calibrated by $(E_{inc},\vec{p}_{inc}/|\vec{p}_{inc}|)$ , the incident energy and normalised momentum direction of the incident photon. The decision to use Cartesian coordinates (rather than angles) is detailed in section 3.2.
@@
The reductions from the design of ShowerFlow in CaloClouds2 (see appendix A) remove redundancies that created submanifolds in the normalising flow. The flow interleaves 12 affine couplings and 2 spline couplings. This has been reduced from the 60 affine couplings and 10 spline couplings used in CaloClouds2, as a much simpler flow was seen to be sufficient to capture the complexity of the distribution, while being faster at inference time and more stable to train. A variant of ShowerFlow that predicted the log of the energy and points per layer was experimented with, but it was found to be unstable, and so not ultimately pursued.
The CoG of the shower is no longer predicted by ShowerFlow, the reasoning for which is discussed later in this section.
The ShowerFlow model is trained using an Adam [52] optimiser, with a learning rate of $0.0002$. The batch size is $62$ and a linear scheduler is used. Due to the high volume of training data, overfitting is not observed.
This output is reverted to physical values in the Flow Calibration, which yields the number of points in each layer, and the energy in each layer.
Then, like in CaloClouds2, a diffusion model which has been distilled to a single step [^4] As with CaloClouds2, distillation of the diffusion model is seen to have no adverse impact on the physics performance of the model. is used to generate individual points as an iid distribution. Unlike CaloClouds2, the diffusion model is no longer conditioned on the total number of points. Removing this conditioning was seen to have no impact on the physics performance of the model; the conditioning is the same as for ShowerFlow (incident particle energy and direction). Raw points express their energy on a log scale, such that the energy spectrum of the points covers all real numbers. These raw points are calibrated in the Final Calibration. Starting from the first layer the photon reaches and working outwards:
@@
    The number of points predicted by ShowerFlow for this layer are selected from the unused points generated from the diffusion model, starting with those generated at the base of the shower. The selected points are moved along the shower axis to sit at the centre of the layer.
@@
On a basic level, detector readout can be mimicked by converting the energy deposits back to detector coordinates and allocating them to cells using records of the cell and layer boundaries. Energy in each cell is summed, and energy that falls into gaps between sensitive material is abandoned. This is done when comparing CaloClouds2 to CaloClouds3, as CaloClouds2 does not have the angular conditioning capability required for use in full reconstruction.
When working with only CaloClouds3 a more extensive framework can also facilitate reconstruction. The DD4hep [53] toolkit exists to provide detector descriptions. Built on top of this, the DDsim and DDML [54] libraries collectively run simulations. A full Geant4 simulation can be run, or an augmented Geant4 simulation, in which particles identified as suitable for a fast simulation model will be delegated to that fast simulation.
@@
Removal of Centre of Gravity (CoG) calibration.
Typically, the CoG is close to the incident particle axis i.e. the shower axis. When the CoG does deviate significantly from the shower axis, it has been shifted by high energy secondary particles scattered at wide angles. As the diffusion model draws points iid, these correlated structures cannot be replicated, and so the tails of the CoG distribution are not well modelled by the diffusion model.
@@ 4.1 Impact of the changes from CaloClouds2 to CaloClouds3
The fidelity of photons with orthogonal impact, between $10$ and $90~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, is compared between CaloClouds2, CaloClouds3 and Geant4, since this is the only region in which CaloClouds2 can produce output. The results are projected into a realistic detector map, as described at the end of section 3.3, and key shower observables calculated – these observables are shown in figures 45 and 8.
@@
The central plot of figure 4 shows the radial profile of the energy. One can observe a slight dip at 0 radius because the impact point is in the centre of a cell, so the first bin is only half a cell wide. Overall, both models replicate the Geant4 data well, with CaloClouds2 producing a marginally flatter ratio.
@@
In the centre plot of figure 5, the number of active cells in concentric rings about the shower axis is shown. Towards 0 radius, bins encompass a smaller volume of the detector, creating a dip in the number of active cells (hence this isn’t a true occupancy).In this metric, both models are excellent, with CaloClouds3 recreating the centre of the distribution the best.
The rightmost of the three plots shows the number of active cells per layer (or the occupancy in each layer). The ability of the models to replicate this pattern is defined by the ShowerFlow component, which specifies the points in each layer. In this plot a slight pattern from the alternating absorber placement can be seen, in the same way as seen for the energy in figure 4, figure 5 shows an excess of hits in every other layer. The change in absorber thickness at layer $20$ is also somewhat visible here. Both models replicate these effects well, but CaloClouds3 has a flatter ratio to Geant4.
@@
CaloClouds2 CaloClouds3
@@
Since the changes from CaloClouds2 to CaloClouds3 are significant, involving both a reduction in size, and the extension of the model to cover incident particles of all angles, It is satisfactory that the metrics in table 2 indicate comparable performance. Overall it appears that the upgrades to CaloClouds3 have not adversely impacted the performance of the model.

### v2

@@ 1 Introduction and Motivation
In this paper, we present the latest of a series of models designed for the fast simulation of photon showers in a high granularity calorimeter, such as those depicted in figure 1. This is the latest of the CaloClouds series of models, which represent the showers to be modeled as pointclouds. Two ML techniques are used; a normalising flow combined with a diffusion model [^2] The code is avaliable at https://github.com/FLC-QU-hep/CaloClouds-3. For a review covering other possible architectures, see [32]. For a comparison of recent models, see the companion paper, a first full physics benchmark for highly granular calorimeter surrogates [16] and results of the CaloChallange [39].
@@
The previous model, CaloClouds2, was only capable of simulating photons incident at a single angle and position. Incident angle conditioning, along with location agnostic training data, allows a single model to be used for all photons entering the calorimeter, which simplifies training, and potentially occupies less memory at inference time. The octagonal barrel shape and magnetic field, which is aligned with the beam axis, break the rotational symmetry about the shower axis, therefore the model must be conditioned on both polar ($\theta$) and azimuthal ($\phi$) directions separately.
@@
In section 2 we describe the example detector we have chosen to demonstrate our architecture on, and the dataset generation is detailed. In section 3 the new CaloClouds3 model is defined. In sections 4 and 5 we compare physics accuracy between the previous iteration, CaloClouds2, and the new model, CaloClouds3. After this, a snapshot of the newly attained capacity for full reconstruction is offered in section 5, however this is covered in more depth in a companion paper [16]. In the penultimate section, section 6, the improvements in inference speed between CaloClouds2 and CaloClouds3 are demonstrated. Finally, we draw the conclusions in section 7.
@@ 2.2 Generation and preprocessing
Photons between $1$ and $127~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ were generated using Geant4, a slightly expanded range from the energies considered in CaloClouds2. $127~$\mathrm{G}\mathrm{e}\mathrm{V}$\$ has been chosen as it is just over half of the initial collision energy planned at the ILD [59]. In the framework used here, a fast model is only utilised (triggered) once a photon in an appropriate energy range reaches the surface of the calorimeter, therefore any photons that start showering before the surface of the calorimeter will not be simulated with the fast simulation model [^3] Potentially, some secondary photons from the shower of an early showering photon could reach the calorimeter surface with sufficient energy to trigger the fast simulation, but this would anyway have the same physics as a low energy primary photon, and so is covered by the dataset.. Consequently it is important to generate the datasets used to train the model without any pollution from early showering photons [^4] For further description of how the model is triggered in the context of a complete event simulation, see section 3.3.. To avoid any early showering, the photons in the training set are inserted into the simulation at the surface of the calorimeter. The angle of each photon is consistent with having originated from the Interaction Point (IP), having $30<\theta<150$ and $80<\phi<110$, sampled to have an even density on the upper barrel segment. Geant4 then records the hits of the shower in the sensitive layers. The training data includes $3\times 10^{6}$ events, with a test set of $7\times 10^{4}$ which was held back.
@@
The centre of each layer is shifted in the $x$ and $y$ directions to remove the shower’s tilt. This shift corresponds to the rotation that would align the incident photon with the $z$ axis, so the size of the shift is calculated from the depth of the layer $d_{\mathrm{layer}}$ and the momentum of the incident particle $\vec{p}_{\mathrm{inc}}$; in the $x$ direction, the shift is $\Delta x=d_{\mathrm{layer}}\tan(\sin^{-1}(\vec{p}_{\mathrm{inc},x}/|\vec{p}_{\mathrm{inc}}|))$, and likewise for the $y$ direction. To first approximation, the shifted showers are now a similar oblong shape, somewhat akin to one created by an upward moving particle. This preprocessing helps the diffusion model to learn the details of backscattering, rather than being forced to focus on getting the gross shape right, which is known from the direction of the incident particle.
Finally, all 30 layers of the detector are trimmed to a box of $100\times 100$ cells, cutting to $\pm 250~$\mathrm{m}\mathrm{m}$\$ about the shower axis. To strike a balance between the fidelity of all Geant4 hits, and computational efficiency, the energy in each layer is binned into a grid of $500\times 500$ bins (each cell is split into $5\times 5$ bins). This is a slightly coarser granularity than that of CaloClouds2, and was implemented because lowering the number of points was seen to have minimal impact on the kinematics after projection, while improving the computational efficiency of the model. As with CaloClouds2, in each event, this grid is given a uniform random offset of up to half a cell, to reduce discretisation effects. Multiple hits may be present in a single bin, and often are near the centre of the shower. Rather than place all the energy of the hits in the bin in a point at the centre of the bin, a point a is created at the location of the highest energy hit in the bin. This slightly better reflects the true energy distribution, at minimal cost. This creates a set of up to $6,000$ points per shower. The changes from the previous version, CaloClouds2, are summarised in table 1.
@@ 3.1 Previous CaloClouds2 model
The key points for the following discussion are that the CaloClouds models are composed of a normalising flow, referred to as ShowerFlow, which determines the overall structure of each photon shower, and a diffusion model, which generates each point. Both models are conditioned to create behaviour appropriate for the incident photon. The previous CaloClouds2 model was only capable of simulating photons at a fixed position, and the model’s hyperparameters were adopted from standard practice for these models.
@@ 3.2 Upgrade methodology
Conditioning on the incident angle is a relatively simple change. In CaloClouds2 the normalising flow was conditioned on $E_{\mathrm{inc}}$, the incident particle energy, in CaloClouds3 it will be conditioned on $(E_{inc},\vec{p}_{inc}/|\vec{p}_{inc}|)$ , the incident energy and normalised momentum direction. Likewise, the diffusion model of CaloClouds3 will be conditioned on $(E_{inc},\vec{p}_{inc}/|\vec{p}_{inc}|)$. Conditioning the diffusion model on the number of points in the shower, used in CaloClouds2, was found to not have a significant effect and therefore removed.
@@
In order to limit the computational resources (and therefore the carbon footprint) required to investigate variations of normalising flow models, some toy models, representing expected data characteristics were employed. These small scale experiments focused on investigating two potential properties of a probability distribution function (pdf); sub-manifolds and first or second order discontinuities. Physics data might be expected to contain sub-manifolds, as the number of output parameters generally exceeds the number of input parameters, and would be expected to be approximately smooth, with few first order discontinuities. For example, the pdf of the energy in each layer of the detector would be expected to vary smoothly with incident particle energy, and have an approximate sub-manifold defined by adjacent layers having similar energies. Toy datasets were created from various pdfs, with and without the aforementioned properties. By observing the training and goodness of fit for normalising flow models with a similar architecture on those toy datasets, we could economically obtain indications of what might be expected in more complex cases. While this is far from conclusive, it helped guide development of the more expensive training process of the full model.
All expected characteristics of the physics data favoured the use of simple normalising flow models. By reducing the parameter space of our ShowerFlow model, it was seen that the number of parameters could be safely reduced by a factor of 5 (but that a factor of 10 was detrimental to performance). In a similar fashion, the number of parameters in the diffusion model was reduced by a factor of 4.
@@
While the toy models showed no overwhelming failure to learn due to manifold overfitting, sub-manifolds did somewhat impair training stability. Identifying and removing sub-manifolds in the normalising flow of our physical model still offered a marginal performance increase. Comparing the training process of ShowerFlow in CaloClouds2 and CaloClouds3 (figure 2), there are indications that a more stable training is taking place in CaloClouds3.
@@ 3.3 CaloClouds3 architecture
Figure 3: The model architecture of CaloClouds3. Parts belonging to ShowerFlow are in purple, and those belonging to the diffusion model are in green. A full description is provided in section 3.3.
As with CaloClouds2, the model uses a normalising flow, referred to as ShowerFlow, (shown in purple in figure 3) and a distilled diffusion model (shown in green in figure 3). ShowerFlow is now limited to distributions along the shower axis, specifically the points and energy per layer, and the diffusion model is still used to create individual points. Both are rescaled by $(E_{inc},\vec{p}_{inc}/|\vec{p}_{inc}|)$ , the incident energy and normalised momentum direction of the incident photon. The decision to use Cartesian coordinates (rather than angles) is detailed in section 3.2.
@@
The reductions from the design of ShowerFlow in CaloClouds2 (see appendix A) remove redundancies that created sub-manifolds in the normalising flow. The flow interleaves 12 affine couplings and 2 spline couplings. This has been reduced from the 60 affine couplings and 10 spline couplings used in CaloClouds2, as a much simpler flow was seen to be sufficient to capture the complexity of the distribution, while being faster at inference time and more stable to train. A variant of ShowerFlow that predicted the log of the energy and points per layer was experimented with, but it was found to be unstable, and so ultimately not pursued. The Center of Gravity (CoG) of the shower is no longer predicted by ShowerFlow, the reasoning for which is discussed later in this section.
The ShowerFlow model is trained using an Adam [38] optimiser, with a learning rate of $0.0002$. The batch size is $62$ and a linear scheduler is used. Due to the high volume of training data, overfitting is not observed. This output will be reverted to physical values in the step Apply macro features, which yields the number of points in each layer, and the energy in each layer.
Then, like in CaloClouds2, a diffusion model which has been distilled to a single step [^5] As with CaloClouds2, distillation of the diffusion model is seen to have no adverse impact on the physics performance of the model. is used to generate a batch of individual points from an independent and identically distributed (iid) distribution. Until the generated points are placed in a layer they are referred to as raw points. Unlike CaloClouds2, the diffusion model is no longer conditioned on the total number of points. Removing this conditioning was seen to have no impact on the physics performance of the model; the conditioning is the same as for ShowerFlow (incident particle energy and direction). Raw points express their energy on a log scale, such that the energy spectrum of the points covers all real numbers. These raw points are arranged as a shower during the step Apply macro features. Starting from the first layer the photon reaches and working outwards:
@@
    The number of points predicted by ShowerFlow for this layer are selected from the raw points generated from the diffusion model, starting with those generated at the base of the shower. These points are moved along the shower axis to sit at the centre of the layer and assigned to that layer, removing them from the raw points.
@@
    The energy of the selected points is returned to a linear scale.
  3. 3.
@@
On a basic level, detector readout can be mimicked by converting the energy deposits back to detector coordinates and allocating them to cells using records of the cell and layer boundaries. To calculate the energy recorded by a cell, the energy of all the points that fall within each cell are summed. Energy that falls into gaps between sensitive cells is abandoned. This is the method used to emulate detector readout with which CaloClouds2 and CaloClouds3 are compared.
When working with only CaloClouds3 a more extensive framework can also facilitate reconstruction. The DD4hep [30] toolkit exists to provide detector descriptions. Built on top of this, the DDsim and DDML [37] libraries collectively run simulations. A full Geant4 simulation can be run, or an augmented Geant4 simulation, in which particles identified as suitable for a fast simulation model will be delegated to that fast simulation. This is only possible with CaloClouds3 as it requires angular conditioning, which CaloClouds2 lacks.
@@
Removal of CoG calibration.
For a perfect detector, conservation of momentum would ensure that the position weighted energy sum of the shower’s energy depositions was exactly on the incident particle axis, also refered to as the shower axis, $\hat{s}$. The ILD has a sampling fraction of around $1\%$ in the ECAL, therefore missing energy may break this symmetry.
In this paper, we define CoG as the weighted sum of the shower’s energy, in a coordinate system relative to the shower axis. Let $c$ be the vector from the point the photon meets the calorimeter to the energy weighted mean of the shower, and $\hat{z}$ be the detector global $z$ direction, that is the direction of the beam then the centre of gravity can be expressed as
$\mathrm{center~of~gravity~X}=((\hat{s}\times\hat{z})\times\hat{s})\cdot c$
and
$\mathrm{center~of~gravity~Y}=(\hat{s}\times\hat{z})\cdot c.$
Typically, both these values are small, because an even sampling of the random energy deposits means the CoG remains close to the shower axis. When the CoG does deviate significantly from the shower axis, it has been shifted by high energy secondary particles scattered at wide angles. As the diffusion model draws points iid, these correlated structures cannot be replicated, and so the tails of the CoG distribution are not well modelled by the diffusion model.
@@
We also consider a third axis for the CoG, in the direction perpendicular to the layers of the detector. This is referred to as center of gravity Z, and is determined by the energy per layer, as modelled by ShowerFlow.
@@ 4.1 Impact of the changes from CaloClouds2 to CaloClouds3
The fidelity of photons with orthogonal impact, between $10$ and $90~$\mathrm{G}\mathrm{e}\mathrm{V}$\$, is compared between CaloClouds2, CaloClouds3 and Geant4, since this is the only region in which CaloClouds2 can produce output. The results are projected into a realistic detector map, as described at the end of section 3.3, and key shower observables calculated – these observables are shown in figures 45 and 6.
@@
The central plot of figure 4 shows the radial profile of the energy. Detector cells are assigned to a bin of this histogram according to their radial distance from the shower axis, and the energy from that cell added to the bin total. One can observe a slight dip at 0 radius because the impact point is in the centre of a cell, so the first bin is only half a cell wide. Overall, both models replicate the Geant4 data well, with CaloClouds2 producing a marginally flatter ratio, and CaloClouds3 tending to overestimate the amount of energy in the tails. We do not know why CaloClouds3 overestimates the tails of the radial distribution.
@@
In the centre plot of figure 5, the number of active cells per unit radius is shown. Like the radial energy, detector cells are assigned to a bin of this histogram according to their radial distance from the shower axis, and the number of cells with energy above half a MIP is counted towards the bin total. Towards 0 radius, bins encompass a smaller volume of the detector, creating a dip in the number of active cells. In this metric, both models are excellent, with CaloClouds3 recreating the centre of the distribution the best.
The rightmost of the three plots shows the number of active cells per layer (or the occupancy in each layer). The ability of the models to replicate this pattern is defined by the ShowerFlow component, which specifies the points in each layer. Due to the alternating absorber placement, every other layer of the calorimeter tends to accumulate more hits. This can be seen in figure 4 as an excess of energy in every other layer, and in figure 5 as an excess of hits in every other layer. The change in absorber thickness at layer $20$ is also somewhat visible here. Both models replicate these effects well, but CaloClouds3 has a flatter ratio to Geant4.
@@
Variable Jenson Shannon divergence
compared CaloClouds2 CaloClouds3
@@
Since the changes from CaloClouds2 to CaloClouds3 are significant, involving both a reduction in size, and the extension of the model to cover incident particles of all angles, it is satisfactory that the metrics in table 2 indicate comparable global performance. The changes have improved some fits, and degraded others, but for the most part CaloClouds3 remains at the same order of magnitude or better than CaloClouds2. Only for the centre of gravity in X/Y does CaloClouds2 perform a lot better, and this is due to the removal of CoG calibration, which is discussed in section 3.3. Overall it appears that the upgrades to CaloClouds3 have not adversely impacted the performance of the model.